OpenStudy (anonymous):

Solve the equation algebraically 5rad(x+4) =absolute value (x-3)

OpenStudy (anonymous):

\[5\sqrt{x+4}=\left| x-3 \right|\]

OpenStudy (jdoe0001):

\(\large { 5\sqrt{x+4}=| x-3 |\\\quad\\ \implies \begin{cases} +(5\sqrt{x+4})=x-3\implies 5\sqrt{x+4}=x-3\\ 5\sqrt{x+4}+3=x \\\hline\\ -(5\sqrt{x+4})=x-3\implies -5\sqrt{x+4}=x-3\\ -5\sqrt{x+4}+3=x \end{cases} }\) two cases, for an absolute value expression, thus two values for "x" keep in mind you always have a + and - case with absolute value expressions

OpenStudy (anonymous):

ok i got that but I still dont know how to solve this

OpenStudy (jdoe0001):

ohh shoot... aehmm yeah... ok... we do have an "x" on the left side... I see

OpenStudy (jdoe0001):

have you ... covered the factoring quadratics?

OpenStudy (anonymous):

idk maybe

OpenStudy (jdoe0001):

hmmm meaning, have you factored something like \(x^2+3x-4\) yet?

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

ok.. have you covered the quadratic formula yet?

OpenStudy (anonymous):

yass

OpenStudy (jdoe0001):

k

OpenStudy (jdoe0001):

let us pick the 1st case, one sec

OpenStudy (jdoe0001):

\(\bf 5\sqrt{x+4}=x-3\impliedby \textit{let us square both sides} \\ \quad \\ (5\sqrt{x+4})^2=(x-3)^2\implies (5)^2(\sqrt{x+4})^2=x^2-6x+9 \\ \quad \\ 25(x+4)=x^2-6x+9\implies 25x+100=x^2-6x+9 \\ \quad \\ 0=x^2-31x-91\) now, we end up with a quadratic equation 91 is a prime, pretty sure of that so it won't factor nicely, so we could get the middle term of -6 so you'd need to use the quadratic formula for that

OpenStudy (jdoe0001):

\(\bf 0=x^2-31x-91\qquad \qquad \textit{quadratic formula} \\ \quad \\ 0={\color{blue}{ 1}}x^2{\color{red}{ -31}}x{\color{green}{ -91}} \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (jdoe0001):

now let use pick the 2nd case \(\bf -5\sqrt{x+4}=x-3\impliedby \textit{let us square both sides} \\ \quad \\ (-5\sqrt{x+4})^2=(x-3)^2\implies (-5)^2(\sqrt{x+4})^2=x^2-6x+9 \\ \quad \\ 25(x+4)=x^2-6x+9\implies 25x+100=x^2-6x+9 \\ \quad \\ 0=x^2-31x-91\) notice, the -5, turns to a positive 25 and we end up with the same equation as the 1st case so, using the quadratic formula for the 1st one, solves also the 2nd one

OpenStudy (anonymous):

ok thank you sooooo much @jdoe0001

OpenStudy (jdoe0001):

ywq

OpenStudy (jdoe0001):

yw rather